Physics Help? We all have had the experience that a room full of people always feels warmer than when the room is empty...
media%2Fef8%2Fef8494d2-4f11-41f9-a9bb-bf
Answer #1

Rate of Heat given off by each person is, R= 100W

Then Heat given by each person in 10 min or (10 * 60) seconds is

   Q= R \times t(seconds)= 100 \times 10 \times 60= 60000J

And Total Heat given off by 10 people in the room is

  Q_t= Q \times 10 = 60000 \times 10= 6 \times 10^5J

Now Volume of the room is V= (4.0 \times 6.0 \times 3.0)m^3= 72m^3

Density of the air, \rho= 1.2 kg/m^3

Then Mass of air is M= Volume \times \rho= 72 \times 1.2= 86.4 kg

Given Specific Heat at constant pressure is c_p= 1005 J/kg. ^{\circ} C

Then Using the fornula,

Heat Gain, Q_t= Mc_p(T_{final}- T_{initial})

Using given initial value of temperature and all other calculated values in above, we get

  6 \times 10^6= (86.4)(1005)(T_{final}- 20)

T_{final}- 20= 6.91

or   T_{final}= 26.91 ^{\circ} C(ANS)